OpenStudy (anonymous):

Let h = (v o^2 / 4.9) sin theta cos theta model the horizontal distance in meters traveled by a projectile. If the initial velocity is 52 meters/second, which equation would you use to find the angle needed to travel 200 meters? A. 551.84sin(2theta)=200 B. 275.92sin(2theta)=200 C. 10.61 sin(2theta)=100 D. 200 sin(2theta)=200

OpenStudy (anonymous):

I do not know where to go from this stage: \[551.84\sin \Theta\]

OpenStudy (kropot72):

You should use the double angle formula: \[\large \sin 2\theta=2 \sin\theta \cos\theta\]

OpenStudy (anonymous):

just tell me the proper equation

OpenStudy (kropot72):

Then we get \[\large 200=\frac{52^{2}}{4.9}\times \frac{\sin 2\theta}{2}\]

OpenStudy (anonymous):

Oh I understand! So that would get me the answer to be B right?

OpenStudy (kropot72):

Correct.

OpenStudy (anonymous):

Great! Thank you! It's always setting up the equation that gets me.

OpenStudy (kropot72):

np. You're welcome :)